Solve the differential equation
dy/dx=- 2xy²
1
Expert's answer
2020-03-17T11:07:31-0400

We use the variable separable method for this question

dy/dx=-2xy2

dy/y2=-2xdx

Integrating both sides, we get

-1/y+c=-x2

1/y=x2+c

Thus, y(x)=1/(x2+c) is the general solution to the differential equation